# Non-perturbative determination of the sphaleron rate for first-order phase transitions

### v1

@@ V Beyond the Standard Model phase transitions
This general strategy has been previously discussed or (at least partly) applied in Kajantie et al. (1996b); Cline and Kainulainen (1996); Losada (1997); Laine (1996); Moore (1999); Brauner et al. (2017); Andersen et al. (2018); Gorda et al. (2019); Niemi et al. (2019); Gould et al. (2019); Friedrich et al. (2022); Ramsey-Musolf et al. (2024); Li et al. (2025).
@@
Hence, by applying the above strategy one can extract reliable partly non-perturbative information about the phase structure of any model without resorting to more conventional (and purely perturbative) approaches, such as direct numerical minimisation of a thermal effective potential which is numerically more expensive compared to working with the $\{x(T),y(T)\}$-mapping, and unreliable for weak transitions unable to find a boundary between first-order transitions and crossovers. On the other hand, one still has to ensure that the perturbative part of the analysis, the dimensional reduction mapping, is reliable including the omission of marginal operators.666 For recent studies where marginal operators cannot be neglected, see Laine et al. (2018); Chala et al. (2024); Bernardo et al. (2025); Chala and Guedes (2025). We discuss this aspect further below.
@@
The strategy outlined above is general for all one-step phase transitions irrespective of a representation of $S$ or multiple such fields, and corresponding dimensional reduction mappings can be obtained in automated fashion Ekstedt et al. (2023). Before moving on, we emphasize that this strategy cannot be used to study multistep phase transitions with more complicated phase transition pattern, see e.g. Weinberg (1974); Land and Carlson (1992); Hammerschmitt et al. (1994); Patel and Ramsey-Musolf (2013); Blinov et al. (2015); Niemi et al. (2021a); Gould and Tenkanen (2024), yet such transitions are located somewhere in-between boundaries (i) and (ii) that can themselves be found, as already emphasized, simply by zero temperature physics.777 Possible existence of multistep transitions in this region is intuitive in the following sense: in the region where the electroweak vacuum is not global, the system has to transition from the confinement phase to the $S$-phase at some high temperature, and only in a vicinity of this region it is possible that system transitions first to the $S$-phase at some temperature, and subsequently to the electroweak phase at some lower temperature. Where exactly this happens in-between boundaries (i) and (ii), requires to study thermal EFT with new scalars actively included Niemi et al. (2021a); Gould and Tenkanen (2024).
@@
Figure 6: Regions of crossover in yellow ($x_{c}>0.1$), weak first-order transitions in orange ($0.10.025<x_{c}<$ 0.1), strong first-order transitions in red ($0.0250<x_{c}<$ 0.025) and regime where marginal operators or other higher-order effects are non-negligible in blue ($x_{c}<0$) in $(M_{S},a_{2})$-plane with fixed $b_{4}=1$. In first column, soft contributions from the singlet are computed at one-loop, and in the second column at two loops, while different rows use $N=1,3,5$, in descending order.
@@
We assume this thermal mass to be heavy compared to the scale of the transition, i.e. $\mu^{2}_{S,3}\sim m^{2}_{D}\sim(gT)^{2}$, and hence integrate out the singlet (akin to temporal gauge field modes).888 One could also integrate the singlet out together with non-zero Matsubara modes in the dimensional reduction, see Laine and Losada (2000); Brauner et al. (2017), yet here we do not discuss this option, for simplicity. Consequently, Higgs thermal self-coupling at the scale of the transition reads
@@
where $a_{2,3}=T\Big{(}a_{2}-\beta_{a_{2}}L_{b}\Big{)}$, and we have included singlet contributions from the hard scale $\sim\pi T$ at one-loop, and from the scale $\sim gT$ (“soft” scale) at two-loop level. [^9] Contributions from the scale $gT$ are enhanced, and reaching the same formal accuracy as for the hard scale contributions requires a higher loop order calculation Kajantie et al. (1996c). Here $L_{b}\propto\ln(\Lambda/T)$ in the hard scale part, and $\beta$ denotes one-loop beta functions describing running with respect to the (parent theory) renormalisation scale $\Lambda$.
@@
where we isolated the SM part (see e.g. Kajantie et al. (1996c); Niemi et al. (2021b)), $b_{4,3}=T(b_{4}-\beta_{b_{4}}L_{b})$ Niemi et al. (2021b) and $\Lambda_{3}$ is the renormalisation scale of the 3D EFT. We note that here the second term describes hard scale corrections at one-loop, but we did not write down the corresponding two-loop contribution from the hard scale, see Niemi et al. (2021b). In practical calculations, it is crucial to include such two-loop contribution to cancel the running of the one-loop term Arnold and Espinosa (1993); Gould and Tenkanen (2021). In Eq. (V) the last two terms are soft scale corrections from the singlet, at one and two loops, respectively.
We note, that for generic models the dimensional reduction matching relations (as well as required one-loop $\beta$-functions) can be obtained using DRalgo code Ekstedt et al. (2023); Fonseca (2021). However, zero-temperature vacuum renormalisation procedure required for relating Lagrangian parameters of a model to physical quantities such as pole masses (see e.g. Kajantie et al. (1996c); Laine et al. (2017); Niemi et al. (2021b)) is not included. Likewise, the soft scale corrections to scalar couplings (c.f. Eq. (18)) is not (automatically) included in the current version.101010 Here we concretely refer to DRalgo version 1.3. However, two-loop soft corrections to the scalar coupling (Eq. (18)) can be obtained (manually) with a following trick: DRalgo can be used to compute the two-loop effective potential $V^{\text{3D}}_{\text{eff}}(v_{3})$ with respect to the Higgs background field $v_{3}$, which includes soft scale corrections by a new scalar $S$. Since the effective potential is the generator of $n$-point Green’s functions $G_{n}$, one can expand it as series in $v_{3}$, identify and isolate contributions of $S$ and extract the required 4-point function from $\frac{1}{4}v^{4}_{3}G_{4}$ for the two-loop matching Kajantie et al. (1996c).
@@
As an illustration of the previous discussions we fix $b_{4}=1$ and solve $T_{c}$ and $x_{c}$ in parts of $(M_{S},a_{2})$-plane where $\mu^{2}_{S}>0$ (and hence also the electroweak minimum is automatically global). We depict result in Fig. 6 which shows different regions of crossover ($x_{c}>0.1$), weak first-order transitions ($0.10.025<x_{c}<$ 0.1), strong first-order transitions ($0.0250<x_{c}<$ 0.025) and regions where marginal operators or other higher-order effects are non-negligible ($x_{c}<0$). In the left column we show results including only one-loop corrections of the singlet from the soft scale, while the right column also includes the two-loop (soft scale) contributions. For each row we use $N=1,3,5$ in descending order.
In top row with $N=1$ we observe that integrating out the singlet at one-loop (left) results in a narrow band of strong transitions with $0.0250<x_{c}<$ 0.025, yet inclusion of two-loop corrections (right) completely removes this region, and even in the top right corner with large $a_{2}$ and $M_{S}$, $x_{c}$ does not become small enough for strong transitions. This result indicates that for $N=1$ there exists no strong one-step phase transitions which can be described within our approach. Yet we emphasize that our analysis does not indicate that this model with $N=1$ could not admit strong two-step transitions in which a light singlet plays a dynamical role and cannot be integrated out.
@@
Figure 8: On the left panel the orange curve shows the $\{x(T),y(T)\}$-mapping for a non-$Z_{2}$ symmetric $N=1$ scalar, with $M_{S}=400~{}$GeV, $a_{2}=3.17$, $\sin\theta=0.1$, $b_{4}=1$, $b_{3}=0$, and two-loop corrections; as in the middle row of Fig. 7. The shown $\{x(T),y(T)\}$-mapping corresponds to temperature range of $155.9-158.6~{}$GeV. On the right panel we used the interpolated lattice data to get the sphaleron rate for the $\{x(T),y(T)\}$-mapping of the specific model, giving us the sphaleron rate as a function of temperature. The gray standard error bounds are again from the continuum limit error estimate. The horizontal dotted line shows the confinement phase sphaleron rate in the SM case as found in Annala and Rummukainen (2023), here scaled to the dimensionless $\tilde{\Gamma}$. The gray dot-dashed line on the right plot shows the Hubble rate terms in (12) without approximating them as constants.
@@
In Fig. 8 we illustrate results for a single representative parameter point (defined in the plot legend), that maps into $x_{c}<0.025$ (left), and allows us to extract the sphaleron rate as function of temperature (right). We observe, that indeed the temperature dependence in $x(T)$ is weak in the range from $y_{c}$ to $y_{f}$ and $y_{p}$, and that $T_{c}$, $T_{f}$ and $T_{p}$ are all withing 1 GeV around 158 GeV, indicating very minor supercooling down from $T_{c}$. This is a typical trend for these heavy (thermal) portal scalar scenarios.
@@ VI Discussion
Combining this information allowed us to confirm that $y_{p}(x)$ and $y_{f}(x)$ curves intersect, and hence find a condition that sphalerons in the Higgs phase are shut off before the percolation occurs, and this happens when $x_{c}\lesssim 0.025$, assuming that $x(T)$ is only mildly changing with temperature in-between $y_{c}=y(T_{c})$ and $y_{p}=y(T_{p})$.121212 Should this assumption be relaxed, then the more general condition to prevent the sphaleron washout is $x_{f}=x(T_{f})<0.025$ and hence $y_{f}>y_{p}$, i.e. $T_{f}>T_{p}$. Under this assumption, the condition $x_{c}\lesssim 0.025$ provides a necessary,131313 Strictly speaking, the condition $x_{c}\lesssim 0.025$ guarantees that sphalerons can not wash out any baryon asymmetry in the Higgs phase, while at slightly larger values of $x_{c}$ it could be still possible to survive from the double exponential washout, if the initial asymmetry generated in front of the bubble is huge compared to experimentally observed value. For details, see Li et al. (2025). yet not sufficient condition for viability of the electroweak baryogenesis. This condition is slightly stricter than previously obtained by analogous investigations Kajantie et al. (1996b); Moore (1999), and hence our results indicate that to preserve the baryon asymmetry the electroweak phase transition has to be slightly stronger than previously thought.
@@
We can also translate the above upper bound on $x_{c}$ to a lower bound on the discontinuity of the Higgs condensate at the critical temperature (or $y_{c}$). Using $x_{c}=0.025$ results to a lower bound $v/T_{c}\equiv\sqrt{2\Delta\langle\phi^{\dagger}\phi\rangle_{c}g^{2}_{3}}/T_{c}\gtrsim 1.33$ (by using the results from Gould et al. (2022) to obtain the discontinuity). [^14] Note, here we have written the condensate in units of $g_{3}^{2}$ which is what is naturally obtained from the lattice. In the abstract the condensate is directly in 4D units. Here we also used the SM value $g_{3}^{2}\simeq 0.4T$. By computing the (gauge-invariant) condensate in perturbation theory for $x=0.025$, we obtain $v/T_{c}\gtrsim 1.01$ at leading order, and $v/T_{c}\gtrsim 1.36$ at $\text{N}^{4}\text{LO}$ using results of Ekstedt et al. (2022, 2024), with latter providing a fair agreement with the lattice.
@@
To give flavor how our findings can be used to study electroweak phase transitions in a class of BSM theories under these assumptions, we also reviewed a general procedure and exemplified it in terms of a simplified setting with singlet scalar portal. In particular, we demonstrated how the most simple quartic portal is sensitive to higher-order perturbative corrections, yet argued that this feature can be mitigated if such portal comprises of multiple fields or fields that are in higher dimensional representations of electroweak multiplets. Furthermore, we showcased the attractive nature of a possible cubic portal to trigger strong electroweak phase transition, a feature that can be realised in e.g. the non-$Z_{2}$ symmetric singlet-extended SM, for which phenomenological signatures are actively predicted and searched for in the present and future collider experiments, c.f. e.g. Niemi et al. (2024); Antusch et al. (2025) and references therein.
@@ Appendix A Temporal gauge field components
In the effective theory of Eq. (2) temporal gauge field components $A^{a}_{0}$ have been integrated out perturbatively, and their effect is encoded in the parameters of the EFT Kajantie et al. (1996c). This procedure relies upon an approximation that the Debye mass $m_{D}\approx N_{m_{D}}g^{2}T$ of $A^{a}_{0}$ fields dominates over the field-dependent mass generated by the Higgs mechanism. For very strong transitions (at very small $x$), that involve large jump in the Higgs condensate, this perturbative approximation becomes unreliable. Technically, the previous criteria amounts to $m^{2}_{D}\gg h_{3}v^{2}_{3}$, where $h_{3}\approx\frac{1}{4}g^{2}T$ is the coupling of $A^{a}_{0}$ to Higgs field, and $v_{3}$ is the scalar background field (used to generate correlation functions in the EFT matching). For small values of $x$, in the Higgs phase we can estimate the size of the background field by the leading order perturbative result $v_{3}\approx\frac{g_{3}}{8\pi x}$ at the critical temperature Ekstedt et al. (2022), and hence we arrive to a lower bound
@@
for the SM values $g^{2}\approx 0.4$ and $N_{m_{D}}\approx\frac{11}{6}$. While this lower bound is indeed smaller than our condition $\bar{x}\approx 0.025$ for the baryon number preservation in Eq. (13), in the future it could be interesting to revisit our simulations without integrating out the temporal gauge field components (c.f. Kajantie et al. (1993); Farakos et al. (1994a); Kajantie et al. (1996b)), in order to scrutinize the accuracy of our present results. We also note, that in BSM theories where a new portal scalar is charged under SU(2), it would increase the value of $N_{m_{D}}$ (c.f. Gorda et al. (2019); Niemi et al. (2019)) and hence integrating out the temporal gauge fields becomes more reliable.
Additionally, as pointed out in Li et al. (2025), previous simulations D’Onofrio et al. (2012, 2014a); Annala and Rummukainen (2023) have not included the perturbative effect of temporal gluon fields to the Higgs thermal mass (see also Brauner et al. (2017); Schicho et al. (2021)). This effect scales as $\sim g^{2}_{Y}g^{3}_{s}/\pi^{3}$ and appears formally at higher order than NLO in the dimensional reduction Kajantie et al. (1996c). Yet, it has been demonstrated to have a small, but not entirely negligible, effect on thermodynamics and the sphaleron rate Li et al. (2025), due to relatively large numerical values of the top quark Yukawa coupling $g_{Y}$ and the strong coupling $g_{s}$.

### v2

@@ V Beyond the Standard Model phase transitions
This general strategy has been previously discussed or (at least partly) applied in Kajantie et al. (1996c); Cline and Kainulainen (1996); Losada (1997); Laine (1996); Moore (1999); Brauner et al. (2017); Andersen et al. (2018); Gorda et al. (2019); Niemi et al. (2019); Gould et al. (2019); Friedrich et al. (2022); Ramsey-Musolf et al. (2024); Li et al. (2025).
@@
Hence, by applying the above strategy one can extract reliable partly non-perturbative information about the phase structure of any model without resorting to more conventional (and purely perturbative) approaches, such as direct numerical minimisation of a thermal effective potential which is numerically more expensive compared to working with the $\{x(T),y(T)\}$-mapping, and unreliable for weak transitions unable to find a boundary between first-order transitions and crossovers. On the other hand, one still has to ensure that the perturbative part of the analysis, the dimensional reduction mapping, is reliable including the omission of marginal operators. [^8] For recent studies where marginal operators cannot be neglected, see Laine et al. (2018); Chala et al. (2024); Bernardo et al. (2025); Chala and Guedes (2025). We discuss this aspect further below.
@@
The strategy outlined above is general for all one-step phase transitions irrespective of a representation of $S$ or multiple such fields, and corresponding dimensional reduction mappings can be obtained in automated fashion Ekstedt et al. (2023). Before moving on, we emphasize that this strategy cannot be used to study multistep phase transitions with more complicated phase transition pattern, see e.g. Weinberg (1974); Land and Carlson (1992); Hammerschmitt et al. (1994); Patel and Ramsey-Musolf (2013); Blinov et al. (2015); Niemi et al. (2021a); Gould and Tenkanen (2024), yet such transitions are located somewhere in-between boundaries (i) and (ii) that can themselves be found, as already emphasized, simply by zero temperature physics. [^9] Possible existence of multistep transitions in this region is intuitive in the following sense: in the region where the electroweak vacuum is not global, the system has to transition from the confinement phase to the $S$-phase at some high temperature, and only in a vicinity of this region it is possible that system transitions first to the $S$-phase at some temperature, and subsequently to the electroweak phase at some lower temperature. Where exactly this happens in-between boundaries (i) and (ii), requires to study thermal EFT with new scalars actively included Niemi et al. (2021a); Gould and Tenkanen (2024).
@@
Figure 6: Regions of crossover in yellow ($x_{c}>0.1$), weak first-order transitions in orange ($0.025<x_{c}<0.1$), strong first-order transitions in red ($0<x_{c}<0.025$) and regime where marginal operators or other higher-order effects are non-negligible in blue ($x_{c}<0$) in $(M_{S},a_{2})$-plane with fixed $b_{4}=1$. In first column, soft contributions from the singlet are computed at one-loop, and in the second column at two loops, while different rows use $N=1,3,5$, in descending order.
@@
We assume this thermal mass to be heavy compared to the scale of the transition, i.e. $\mu^{2}_{S,3}\sim m^{2}_{D}\sim(gT)^{2}$, and hence integrate out the singlet (akin to temporal gauge field modes). [^10] If the singlet vacuum mass parameter is very heavy ($\mu^{2}_{S}\sim(\pi T)^{2}$), the singlet needs to be integrated out entirely (including the zero Matsubara mode) in the dimensional reduction, see Laine and Losada (2000); Brauner et al. (2017). We do not discuss this situation here for simplicity, c.f. Niemi et al. (2019). The scale of the transition is set by the Higgs thermal mass $m^{2}_{3}=y(T)g^{4}_{3}\sim\frac{1}{x}g^{4}T^{2}$ around the critical temperature $y_{c}\sim 1/x$. For small $x\sim g/\pi$ one hence has $m^{2}_{3}\sim(g^{\frac{3}{2}}T/\sqrt{\pi})^{2}$, which is parametrically in-between the soft $(gT)$ and ultrasoft $(g^{2}T/\pi)$ scales Gynther and Vepsalainen (2006); Ekstedt et al. (2022); Gould and Tenkanen (2024).
Consequently, Higgs thermal self-coupling at the scale of the transition reads
@@
where $a_{2,3}=T\Big(a_{2}-\beta_{a_{2}}L_{b}\Big)$, and we have included singlet contributions from the hard scale $\sim\pi T$ at one-loop, and from the soft scale $\sim gT$ at two-loop level. [^11] Contributions from the scale $gT$ are enhanced, and reaching the same formal accuracy as for the hard scale contributions requires a higher loop order calculation Kajantie et al. (1996a). Here $L_{b}\propto\ln(\Lambda/T)$ in the hard scale part, and $\beta$ denotes one-loop beta functions describing running with respect to the (parent theory) renormalisation scale $\Lambda$.
@@
where we isolated the SM part (see e.g. Kajantie et al. (1996a); Niemi et al. (2021b)), $b_{4,3}=T(b_{4}-\beta_{b_{4}}L_{b})$ Niemi et al. (2021b) and $\Lambda_{3}$ is the renormalisation scale of the 3D EFT. We note that here the second term describes hard scale corrections at one-loop, but we did not write down the corresponding two-loop contribution from the hard scale, see Niemi et al. (2021b). In practical calculations, it is crucial to include such two-loop contribution to cancel the running of the one-loop term Arnold and Espinosa (1993); Gould and Tenkanen (2021). In Eq. (V) the last two terms are soft scale corrections from the singlet, at one and two loops, respectively.
We note, that for generic models the dimensional reduction matching relations (as well as required one-loop $\beta$-functions) can be obtained using DRalgo code Ekstedt et al. (2023); Fonseca (2021). However, zero-temperature vacuum renormalisation procedure required for relating Lagrangian parameters of a model to physical quantities such as pole masses (see e.g. Kajantie et al. (1996a); Laine et al. (2017); Niemi et al. (2021b)) is not included. Likewise, the soft scale corrections to scalar couplings (c.f. Eq. (18)) is not (automatically) included in the current version. [^12] Here we concretely refer to DRalgo version 1.3. However, two-loop soft corrections to the scalar coupling (Eq. (18)) can be obtained (manually) with a following trick: DRalgo can be used to compute the two-loop effective potential $V^{\text{3D}}_{\text{eff}}(v_{3})$ with respect to the Higgs background field $v_{3}$, which includes soft scale corrections by a new scalar $S$. Since the effective potential is the generator of $n$-point Green’s functions $G_{n}$, one can expand it as series in $v_{3}$, identify and isolate contributions of $S$ and extract the required 4-point function from $\frac{1}{4}v^{4}_{3}G_{4}$ for the two-loop matching Kajantie et al. (1996a).
@@
As an illustration of the previous discussions we fix $b_{4}=1$ and solve $T_{c}$ and $x_{c}$ in parts of $(M_{S},a_{2})$-plane where $\mu^{2}_{S}>0$ (and hence also the electroweak minimum is automatically global). We depict result in Fig. 6 which shows different regions of crossover ($x_{c}>0.1$), weak first-order transitions ($0.025<x_{c}<0.1$), strong first-order transitions ($0<x_{c}<0.025$) and regions where marginal operators or other higher-order effects are non-negligible ($x_{c}<0$). In the left column we show results including only one-loop corrections of the singlet from the soft scale, while the right column also includes the two-loop (soft scale) contributions. For each row we use $N=1,3,5$ in descending order.
In top row with $N=1$ we observe that integrating out the singlet at one-loop (left) results in a narrow band of strong transitions with $0<x_{c}<0.025$, yet inclusion of two-loop corrections (right) completely removes this region, and even in the top right corner with large $a_{2}$ and $M_{S}$, $x_{c}$ does not become small enough for strong transitions. This result indicates that for $N=1$ there exists no strong one-step phase transitions which can be described within our approach. Yet we emphasize that our analysis does not indicate that this model with $N=1$ could not admit strong two-step transitions in which a light singlet plays a dynamical role and cannot be integrated out.
@@
Figure 8: On the left panel the orange curve shows the $\{x(T),y(T)\}$-mapping for a non-$Z_{2}$ symmetric $N=1$ scalar, with $M_{S}=400$~GeV, $a_{2}=3.17$, $\sin\theta=0.1$, $b_{4}=1$, $b_{3}=0$, and two-loop corrections; as in the middle row of Fig. 7. The shown $\{x(T),y(T)\}$-mapping corresponds to temperature range of 155.9158.6155.9-158.6~GeV. On the right panel we used the interpolated lattice data to get the sphaleron rate for the $\{x(T),y(T)\}$-mapping of the specific model, giving us the sphaleron rate as a function of temperature. The gray standard error bounds are again from the continuum limit error estimate. The horizontal dotted line shows the confinement phase sphaleron rate in the SM case as found in Annala and Rummukainen (2023), here scaled to the dimensionless $\tilde{\Gamma}$. The gray dot-dashed line on the right plot shows the Hubble rate terms in (12) without approximating them as constants.
@@
In Fig. 8 we illustrate results for a single representative parameter point (defined in the plot legend), that maps into $x_{c}<0.025$ (left), and allows us to extract the sphaleron rate as function of temperature (right). We observe, that indeed the temperature dependence in $x(T)$ is weak in the range from $y_{c}$ to $y_{f}$ and $y_{p}$, and that $T_{c}$, $T_{f}$ and $T_{p}$ are all within 1 GeV around 158 GeV, indicating very minor supercooling down from $T_{c}$. This is a typical trend for these heavy (thermal) portal scalar scenarios.
@@ VI Discussion
Combining this information allowed us to confirm that $y_{p}(x)$ and $y_{f}(x)$ curves intersect, and hence find a condition that sphalerons in the Higgs phase are shut off before the percolation occurs, and this happens when $x_{c}\lesssim 0.025$, assuming that $x(T)$ is only mildly changing with temperature in-between $y_{c}=y(T_{c})$ and $y_{p}=y(T_{p})$. [^14] Should this assumption be relaxed, then the more general condition to prevent the sphaleron washout is $x_{f}=x(T_{f})<0.025$ and hence $y_{f}>y_{p}$, i.e. $T_{f}>T_{p}$. Under this assumption, the condition $x_{c}\lesssim 0.025$ provides a necessary, [^15] Strictly speaking, the condition $x_{c}\lesssim 0.025$ guarantees that sphalerons cannot wash out any baryon asymmetry in the Higgs phase, while at slightly larger values of $x_{c}$ it could be still possible to survive from the double exponential washout, if the initial asymmetry generated in front of the bubble is huge compared to experimentally observed value. For details, see Li et al. (2025). yet not sufficient condition for viability of the electroweak baryogenesis. This condition is slightly stricter than previously obtained by analogous investigations Kajantie et al. (1996c); Moore (1999), and hence our results indicate that to preserve the baryon asymmetry the electroweak phase transition has to be slightly stronger than previously thought.
@@
We can also translate the above upper bound on $x_{c}$ to a lower bound on the discontinuity of the Higgs condensate $\Delta\langle\phi^{\dagger}\phi\rangle_{c}$ Farakos et al. (1995) defined as the difference of the condensate in the different phases at the critical temperature (or $y_{c}$). For the explicit definition of $\langle\phi^{\dagger}\phi\rangle$ see e.g. Annala and Rummukainen (2023) Eq. (14). Using $x_{c}=0.025$ results to a lower bound $v/T_{c}\equiv\sqrt{2\Delta\langle\phi^{\dagger}\phi\rangle_{c}g^{2}_{3}}/T_{c}\gtrsim 1.33$ (by using the results from Gould et al. (2022) to obtain the discontinuity). [^16] Note, here we have written the condensate in units of $g_{3}^{2}$ which is what is naturally obtained from the lattice. In the abstract the condensate is directly in 4D units. Here we also used the SM value $g_{3}^{2}\simeq 0.4T$. By computing the (gauge-invariant) condensate in perturbation theory for $x=0.025$, we obtain $v/T_{c}\gtrsim 1.01$ at leading order, and $v/T_{c}\gtrsim 1.36$ at $\text{N}^{4}\text{LO}$ using results of Ekstedt et al. (2022, 2024), with latter providing a fair agreement with the lattice.
@@
To give flavor how our findings can be used to study electroweak phase transitions in a class of BSM theories under these assumptions, we also reviewed a general procedure and exemplified it in terms of a simplified setting with singlet scalar portal. In particular, we demonstrated how the most simple quartic portal is sensitive to higher-order perturbative corrections, yet argued that this feature can be mitigated if such portal comprises of multiple fields or fields that are in higher dimensional representations of electroweak multiplets. Furthermore, we showcased the attractive nature of a possible cubic portal to trigger strong electroweak phase transition, a feature that can be realised in e.g. the non-$Z_{2}$ symmetric singlet-extended SM, for which phenomenological signatures are actively predicted and searched for in the present and future collider experiments, c.f. e.g. Niemi et al. (2024); Antusch and others (2025) and references therein.
@@ Appendix A Temporal gauge field components
In the effective theory of Eq. (2) temporal gauge field components $A^{a}_{0}$ have been integrated out perturbatively, and their effect is encoded in the parameters of the EFT Kajantie et al. (1996a). This procedure relies upon an approximation that the Debye mass $m_{D}\approx N_{m_{D}}g^{2}T$ of $A^{a}_{0}$ fields dominates over the field-dependent mass generated by the Higgs mechanism. For very strong transitions (at very small $x$), that involve large jump in the Higgs condensate, this perturbative approximation becomes unreliable. Technically, the previous criteria amounts to $m^{2}_{D}\gg h_{3}v^{2}_{3}$, where $h_{3}\approx\frac{1}{4}g^{2}T$ is the coupling of $A^{a}_{0}$ to Higgs field, and $v_{3}$ is the scalar background field (used to generate correlation functions in the EFT matching). For small values of $x$, in the Higgs phase we can estimate the size of the background field by the leading order perturbative result $v_{3}\approx\frac{g_{3}}{8\pi x}$ at the critical temperature Ekstedt et al. (2022), and hence we arrive to a lower bound
@@
for the SM values $g^{2}\approx 0.4$ and $N_{m_{D}}\approx\frac{11}{6}$. While this lower bound is indeed smaller than our condition $\bar{x}\approx 0.025$ for the baryon number preservation in Eq. (13), in the future it could be interesting to revisit our simulations without integrating out the temporal gauge field components (c.f. Kajantie et al. (1993); Farakos et al. (1994b); Kajantie et al. (1996c)), in order to scrutinize the accuracy of our present results. We also note, that in BSM theories where a new portal scalar is charged under SU(2), it would increase the value of $N_{m_{D}}$ (c.f. Gorda et al. (2019); Niemi et al. (2019)) and hence integrating out the temporal gauge fields becomes more reliable.
Additionally, as pointed out in Li et al. (2025), previous simulations D’Onofrio et al. (2012, 2014b); Annala and Rummukainen (2023) have not included the perturbative effect of temporal gluon fields to the Higgs thermal mass (see also Brauner et al. (2017); Schicho et al. (2021)). This effect scales as $\sim g^{2}_{Y}g^{3}_{s}/\pi^{3}$ and appears formally at higher order than NLO in the dimensional reduction Kajantie et al. (1996a). Yet, it has been demonstrated to have a small, but not entirely negligible, effect on thermodynamics and the sphaleron rate Li et al. (2025), due to relatively large numerical values of the top quark Yukawa coupling $g_{Y}$ and the strong coupling $g_{s}$.